Find the mean of the probability distribution of a random variable X which if 𝑃(𝑋) =




π‘₯+1




20




for X= 1, 2, 3, 4, and 5.

1
Expert's answer
2022-03-24T19:18:44-0400

"EX=\sum_{x=1}^5{x\cdot \frac{x+1}{20}}=\frac{1}{20}\left( 1\cdot 2+2\cdot 3+3\cdot 4+4\cdot 5+5\cdot 6 \right) =3.5"